# A Vision for Trustworthy, Fair, and Efficient Socio-Technical Control using Karma Economies

### v1

@@ 1 Introduction
With these pillars in place, we next proceed with formalizing the problem of allocating repeated resources dynamically in a manner that maximizes long-run Nash welfare. First, in Section 3, we take a centralized view on the problem to provide structural insights on Maximum Long-run Nash Welfare (MLNW) solutions with examples. Second, in Section 4 we formally introduce karma economies and their game-theoretical solution concept of Karma Equilibrium (KE); and show that KE coincide with MLNW under natural conditions. This establishes the core building block to operationalize our vision for trustworthy, fair, and efficient socio-technical control: long-run Nash welfare jointly formalizes fairness and efficiency, and it is maximized in a trustworthy manner at the game-theoretical equilibrium of the karma economy. Section 5 thus lays out a possible future smart city re-imagined with karma. The closed nature of the karma economy allows it to control the scope of fairness and efficiency: it provides unprecedented flexibility to decide at what level(s) long-run Nash welfare should be maximized, and correspondingly, whether different karma economies should be coupled or kept separate. Finally, Section 6 concludes the paper, pointing to several promising directions for future research.
Summary of contributions
To summarize, the main contributions of this paper are:
  • Survey contribution: In developing our pillars in Section 2, as well as placing karma economies in the literature in Section 4.5, we review key concepts and results from economics, in particular social choice theory and bargaining theory, with a focus on how these concepts apply to socio-technical control; and conversely, how control-theoretic notions can address open problems pertaining to dynamic resource allocations.
  • Novel insights: Our results on the structure of MLNW in Section 3, as well as the connection between KE and MLNW in Section 4, are novel. These results, whose proofs follow standard techniques and are included in A, constitute essential building blocks to operationalize our vision for trustworthy, fair, and efficient socio-technical control.
  • Visionary perspective: Finally, by developing a rigorously motivated, economically founded, and actionable vision for future smart cities, we hope that this paper will inspire research on dynamic, non-monetary socio-technical control schemes.
@@ 2 Pillars of Trustworthy, Fair, and Efficient Socio-Technical Control
In this section, we take a critical view of the literature on resource allocation, both from economics and control, with the goal of establishing the foundations of karma economies. Our coverage of the literature is distilled on four key messages or pillars that guide our vision for trustworthy, fair, and efficient socio-technical control.
@@ 2.1 Pillar 1: Single-shot and static solutions are limited
It turns out, however, that there is not much that can be done to tailor the allocations to the private users’ urgency, as the celebrated impossibility theorems of social choice theory assert (Arrow, 1950; Gibbard, 1973; Satterthwaite, 1975; Hylland, 1980). We recall here the theorem due to Hylland (1980) which considers more general cardinal preferences and non-deterministic allocation rules. Stating the theorem requires some preliminary notation. Let ${\mathcal{X}}$ denote a set of allocations with at least three elements, e.g., which of the users enter the priority lanes, and let ${\mathcal{N}}=\{1,\dots,n\}$ be the index set of the users. Each user $i\in{\mathcal{N}}$ has a private individual reward function over the allocations $r_{i}:{\mathcal{X}}\rightarrow{\mathbb{R}}$, $r_{i}\in{\mathcal{R}}_{i}$, and $\bm{r}=\left(r_{i}\right)_{i\in{\mathcal{N}}}\in{\mathcal{R}}=\prod_{i\in{\mathcal{N}}}{\mathcal{R}}_{i}$ is the stacked profile of reward functions. An allocation rule $f:{\mathcal{R}}\rightarrow\Delta({\mathcal{X}})$ takes as input a profile of reported reward functions $\bm{\tilde{r}}=\left(\tilde{r}_{i}\right)_{i\in{\mathcal{N}}}\in{\mathcal{R}}$ and outputs a probability distribution over ${\mathcal{X}}$ according to which an allocation is randomly chosen. An allocation rule $f$ is said to be strategy-proof if under $f$, it is a dominant strategy for each user to report its reward function truthfully. An allocation rule $f$ is said to be ex-post Pareto efficient if for all $\bm{r}\in{\mathcal{R}}$, all allocations in the support of $f(\bm{r})$ are Pareto efficient, i.e., they are not dominated by any allocation that makes no users worse off and at least one user better off according to $\bm{r}$. Notice that Pareto efficiency is the most primitive notion of efficiency one could hope for: in Example 2, for instance, any allocation that fully utilizes the priority lanes is Pareto efficient.
{mdframed}
[backgroundcolor=orange!10]
@@
A random dictatorship can be described as follows: choose a user randomly according to some probability distribution and implement its most preferred allocation; if the user has multiple most preferred allocations, choose a second user randomly and implement its most preferred allocation among those pre-selected by the first user; and so on. The intuition behind Theorem 1 should be clear in light of Example 2: any attempt to tailor the allocation to the private users’ urgency will cause them to report the maximum urgency possible. In fact, the simple random controller is precisely a “random dictatorship” in the sense of Theorem 1. Later, in Pillar 3 (Section 2.3), we discuss how repetition and dynamics enable escaping the impossibility of Theorem 1. But first, we argue that the most popular avenue to escape this impossibility – money – is not always well-suited for socio-technical control.
@@ 2.2 Pillar 2: Money does not always cut it
The internet runs on the Transmission Control Protocol (TCP) and its celebrated additive-increase / multiplicative-decrease (AIMD) distributed real-time congestion controller (Chiu and Jain, 1989). Each internet host follows a standard protocol in which it increases its sending rate additively if no congestion is detected, and decreases the rate multiplicatively when congestion is detected. The network does not accommodate hosts that do not follow the standard protocol and their packets are dropped. When all hosts follow the standard protocol, each of their sending rates is guaranteed to converge to the fair share of the network capacity when congestion occurs [^1] In single-link networks, the fair share is the link capacity divided by the number of hosts. In more general networks, different variants of AIMD lead to different notions of fair share, e.g., proportional (Kelly et al., 1998) or max-min (Mo and Walrand, 2000), but they are essentially similar..
One could envision an alternative controller that makes the internet more efficient by giving priority to more urgent flows. To truthfully reveal the private urgency, a real-time bandwidth pricing scheme could be adopted. This would essentially constitute a form of price discrimination that led to a fierce debate in the United States since the 1990s, when it was suggested to deregulate the internet and let Internet Service Providers (ISPs) charge content hosts differently for available bandwidth (Hahn and Wallsten, 2006). The debate culminated in a strong public sentiment for net-neutrality: protecting the internet’s integrity as a free and open resource that plays a crucial role in democratizing our modern societies (Obama, 2016).
{mdframed}
[backgroundcolor=green!10]
@@
Despite of traffic congestion being a severe societal problem in most of the world’s major cities, leading to loss of time, air pollution, and poor quality of life, only a handful of cities have successfully adopted congestion pricing [^2] At the time of writing, those are Singapore, London, Stockholm, Milan, Gothenburg, Oslo, Manhattan, and several other smaller cities and old town centers.. A typical congestion pricing policy charges drivers entering the city center a time-varying toll aimed at mimicking the real-time optimal price that would keep the city free of congestion. The case of Manhattan, NYC is particularly interesting, as it resisted congestion pricing for almost 73 years despite being one of the most congested cities in the world. Namely, William Vickrey, widely regarded as the father of congestion pricing (Vickrey, 1969), originally conceptualized it for the NYC subway system in 1952, while the Manhattan policy finally came into effect in early 2025. The long debate featured several rejected proposals (e.g., in 2008 and 2015) as well as protests and lawsuits against the policy, due to concerns that it would hurt marginalized communities that rely on driving to Manhattan from the boroughs (Burkett, 2024), small business owners (Trapani, 2023), and others (John Dias, 2023). As a consequence of this controversy, the final policy was implemented in January 2025 at a peak toll price of \$9, instead of the originally proposed \$15, which was already a conservative approximation of the theoretically optimal price to eliminate congestion.
Notice that, other than tolling entry into a whole city, there are other real-world implementations of congestion pricing on smaller scales, notably the highway express lanes in California and other US states that largely resemble our running example (cf. Figure 3). These so-called High Occupancy/Toll (HOT) lanes have faced less controversy than tolling a whole city, or even the whole highway, since users at least have access to the free general purpose lanes (Decorla-Souza, 2008). Nonetheless, practical limitations on how high the toll can be persist, and HOT lanes have been nicknamed “Lexus lanes” due to public perception that they are “built for the rich” (Patterson and Levinson, 2008).
{mdframed}
[backgroundcolor=green!10]
@@ 2.3 Pillar 3: Embrace the dynamics
Example 3 demonstrates that repetition, coupled with dynamics that link the allocations across time, opens up new possibilities with respect to the single-shot and static setting. This is formalized by a recent but already seminal result by Jackson and Sonnenschein (2007). Stating this result requires extending the notation introduced in Section 2.1 to repeated and dynamic settings. A finite set ${\mathcal{X}}\subset{\mathbb{N}}$ is repeatedly allocated from over $T\in{\mathbb{N}}$ time-steps. In each time-step $t\in{\mathcal{T}}=\{1,\dots,T\}$, the private individual reward functions $r_{t,i}$ are drawn independently across time and users from a finite set ~$\tilde{\mathcal{R}}_{i}\subset{\mathcal{R}}_{i}$ according to publicly known probability distributions $\bm{\sigma_{i}}\in\Delta(\tilde{\mathcal{R}}_{i})$, with $\bm{\sigma}$ denoting the corresponding joint distribution. Let $f$ be a single-shot allocation rule as introduced in Section 2.1, and let $\bar{r}_{i}^{f}=\operatorname*{\mathbb{E}}_{\bm{r}\sim\bm{\sigma},x\sim f(\bm{r})}\left[r_{i}(x)\right]$ denote user $i$’s expected single-shot reward under $f$ (if all users report their rewards truthfully). Allocation rule $f$ is said to be ex-ante Pareto efficient if there is no dominating $f^{\prime}$ for which $\bar{r}_{i}^{f^{\prime}}\geq\bar{r}_{i}^{f}$ for all $i\in{\mathcal{N}}$ with strict inequality for at least one $i$. Consider the following linking mechanism aimed at implementing $f$ in a truth-revealing manner when repeated over $T$. Each user $i$ receives a budget to report each $r_{i}\in\tilde{\mathcal{R}}_{i}$ exactly $\sigma_{i}(r_{i})\,T$ times [^4] For the sake of simplicity we assume here that $\sigma_{i}(r_{i})\,T$ is integer for all $i\in{\mathcal{N}}$, $r_{i}\in\tilde{\mathcal{R}}_{i}$; c.f. Jackson and Sonnenschein (2007) for the general treatment., and the allocation in each time-step $t$ is chosen according to $f$ evaluated at the present reports $\bm{\tilde{r}_{t}}$. Let $\bm{\pi_{i}}=\left(\bm{\pi_{t,i}}\right)_{t\in{\mathcal{T}}}$ denote user $i$’s policy under this mechanism that specifies a probability distribution over its reports at each time-step (as a function of its true reward, the history of reports and allocations, etc.), and $\bm{\pi}=\left(\bm{\pi_{i}}\right)_{i\in{\mathcal{N}}}$ the policy profile of all users. When policy profile $\bm{\pi}$ is followed, each user experiences an expected average reward of $\bar{r}_{i}(\bm{\pi})=\cfrac{1}{T}\,\operatorname*{\mathbb{E}}\limits_{\begin{subarray}{c}\bm{\tilde{r}_{t}}\sim\bm{\pi_{t}}\\ x_{t}\sim f(\bm{\tilde{r}_{t}})\\ t\in{\mathcal{T}}\end{subarray}}\left[\sum\limits_{t\in{\mathcal{T}}}r_{t,i}(x_{t})\right]$. A policy profile $\bm{\pi^{\star}}$ is an $\epsilon$-approximate Bayesian Nash equilibrium if no user can unilaterally improve its expected average reward by more than $\epsilon$, i.e., if $\bar{r}_{i}\left(\bm{\pi_{i}^{\star}},\bm{\pi^{\star}_{-i}}\right)\geq\bar{r}_{i}\left(\bm{\pi_{i}},\bm{\pi^{\star}_{-i}}\right)-\epsilon$ for all users $i\in{\mathcal{N}}$ and policies $\bm{\pi_{i}}$. A policy profile in which all users report their rewards as truthfully as possible [^5] In the ex-post realization of rewards, it may not be possible to report them truthfully in every time-step under the linking mechanism, and the reports are as truthful as possible if they are truthful except for a minimal number of time-steps; cf. Jackson and Sonnenschein (2007) for the formal definition. is denoted by $\bm{\pi^{\textup{true}}}$.
{mdframed}
[backgroundcolor=orange!10]
@@
Theorem 2 essentially shows that dynamics enable escaping the static impossibility of jointly Pareto efficient, truthful, and non-dictatorial allocation rules, as established in Theorem 1. To make this claim rigorous, some subtleties should be pointed out between both theorems. The first subtlety is that Theorem 1 concerns ex-post Pareto efficiency while Theorem 2 concerns ex-ante Pareto efficiency; however, the ex-ante notion is stronger and leads to even less possibilities in the single-shot setting (Hylland, 1980, Theorem 2) [^6] Namely, the only possibility to attain ex-ante Pareto efficiency and strategy-proofness in the single-shot setting is a (non-random) dictatorship.. The second subtlety is that Theorem 1 concerns strategy-proofness, i.e., truthful reporting is a dominant strategy; while Theorem 2 concerns so-called Bayesian Nash incentive compatibility, i.e., truthful reporting is an ($\epsilon$-approximate) Bayesian Nash equilibrium, which is weaker than strategy-proofness. However, it is generally believed, though difficult to formally prove, that relaxing strategy-proofness to Bayesian Nash incentive compatibility does not open significant possibilities in the single-shot setting. Some impossibility results for Bayesian Nash incentive compatibility have also been established recently (Ehlers et al., 2020; Kikuchi and Koriyama, 2024).
While Theorem 2 provides hope that trustworthy, fair and efficient dynamic resource allocation schemes are feasible, the linking mechanism used in the theorem is not well-suited for socio-technical control. To summarize the shortcomings of this mechanism in control-theoretic terms, it is ‘open-loop.’ Namely, it requires precise knowledge of the reward function distributions $\bm{\sigma_{i}}$ to issue an exact correct amount of tokens. Moreover, it is not suitable for allocations that are repeated indefinitely, as is the case in most socio-technical control contexts, since it would require issuing infinite tokens. As discussed in Section 4.5, these shortcomings are addressed by our proposed ‘closed-loop’ counter-part, that is, karma economies.
Of these shortcomings, the assumption that the individual reward function distributions are truthfully reported and publicly known is particularly severe for envisioned large-scale socio-technical control applications. While it may not be possible for a dynamic mechanism to truthfully implement all ex-ante Pareto efficient allocation rules without knowing the exact private distributions, it could still be possible for a particular notion of efficiency that does not discriminate between distributions, introduced next.

### v2

@@ 1 Introduction
With these pillars in place, we next proceed with formalizing the problem of allocating repeated resources dynamically in a manner that maximizes long-run Nash welfare. First, in Section 3, we take a centralized view on the problem to provide structural insights on Maximum Long-run Nash Welfare (MLNW) solutions with examples. Second, in Section 4 we formally introduce karma economies and their game-theoretical solution concept of Karma Equilibrium (KE); and show that KE coincide with MLNW under natural conditions. This establishes the core building block to operationalize our vision for trustworthy, fair, and efficient socio-technical control: long-run Nash welfare jointly formalizes fairness and efficiency, and it is maximized in a trustworthy manner at the game-theoretical equilibrium of the karma economy. Section 5 thus lays out a possible future smart city re-imagined with karma. The closed nature of the karma economy allows it to control the scope of fairness and efficiency: it provides unprecedented flexibility to decide at what level(s) long-run Nash welfare should be maximized, and correspondingly, whether different karma economies should be coupled or kept separate. Finally, Section 6 concludes the paper, pointing to several promising directions for future research on dynamic, non-monetary socio-technical control schemes.
@@ 2.1 Pillar 1: Single-shot and static solutions are limited
It turns out, however, that there is not much that can be done to tailor the allocations to the private users’ urgency, as the celebrated impossibility theorems of social choice theory assert (Arrow, 1950; Gibbard, 1973; Satterthwaite, 1975; Hylland, 1980). We recall the theorem due to Hylland (1980) which considers more general cardinal preferences and non-deterministic allocation rules. Stating the theorem requires some preliminary notation. Let $\mathcal{X}$} denote a set of allocations with at least three elements, e.g., which of the users enter the priority lanes, and let ${\mathcal{N}}=\{1,\dots,n\}$ be the index set of the users. Each user $i\in{\mathcal{N}}$ has a private individual reward function over the allocations $r_{i}:{\mathcal{X}}\rightarrow{\mathbb{R}}$, $r_{i}\in{\mathcal{R}}_{i}$, and $\bm{r}=\left(r_{i}\right)_{i\in{\mathcal{N}}}\in{\mathcal{R}}=\prod_{i\in{\mathcal{N}}}{\mathcal{R}}_{i}$ is the stacked profile of reward functions. An allocation rule $f:{\mathcal{R}}\rightarrow\Delta({\mathcal{X}})$ takes as input a profile of reported reward functions $\bm{\tilde{r}}=\left(\tilde{r}_{i}\right)_{i\in{\mathcal{N}}}\in{\mathcal{R}}$ and outputs a probability distribution over $\mathcal{X}$} according to which an allocation is randomly chosen. An allocation rule $f$ is said to be strategy-proof if under $f$, it is a dominant strategy for each user to report its reward function truthfully. An allocation rule $f$ is said to be ex-post Pareto efficient if for all $\bm{r}\in{\mathcal{R}}$, all allocations in the support of $f(\bm{r})$ are Pareto efficient, i.e., they are not dominated by any allocation that makes no users worse off and at least one user better off according to $\bm{r}$. Notice that Pareto efficiency is the most primitive notion of efficiency one could hope for: in Example 2, for instance, any allocation that fully utilizes the priority lanes is Pareto efficient.
@@
A random dictatorship can be described as follows: choose a user randomly according to some probability distribution and implement its most preferred allocation; if the user has multiple most preferred allocations, choose a second user randomly and implement its most preferred allocation among those pre-selected by the first user; and so on. Theorem 1 thus establishes an impossibility of jointly achieving trust and efficiency. The intuition should be clear in light of Example 2: any attempt to tailor the allocation to the private users’ urgency will cause them to report the maximum urgency possible. In fact, the simple random controller is precisely a “random dictatorship” in the sense of Theorem 1.
@@ 2.2 Pillar 2: Money does not always cut it
The internet runs on the Transmission Control Protocol (TCP) and its celebrated additive-increase / multiplicative-decrease (AIMD) distributed real-time congestion controller (Chiu and Jain, 1989). Each internet host follows a standard protocol in which it increases its sending rate additively if no congestion is detected, and decreases the rate multiplicatively when congestion is detected. The network does not accommodate hosts that do not follow the standard protocol and their packets are dropped. When all hosts follow the standard protocol, each of their sending rates is guaranteed to converge to the fair share of the network capacity when congestion occurs [^1] In single-link networks, the fair share is the link capacity divided by the number of hosts. In more general networks, different variants of AIMD lead to different notions of fair share, e.g., proportional (Kelly et al., 1998) or max-min (Mo and Walrand, 2000), but they are essentially similar.. One could envision an alternative controller that makes the internet more efficient by giving priority to more urgent flows. To truthfully reveal the private urgency, a real-time bandwidth pricing scheme could be adopted. This would essentially constitute a form of price discrimination that led to a fierce debate in the United States since the 1990s, when it was suggested to deregulate the internet and let Internet Service Providers (ISPs) charge content hosts differently for available bandwidth (Hahn and Wallsten, 2006). The debate culminated in a strong public sentiment for net-neutrality: protecting the internet’s integrity as a free and open resource that plays a crucial role in democratizing our modern societies (Obama, 2016).
@@
Despite of traffic congestion being a severe societal problem in most of the world’s major cities, leading to loss of time, air pollution, and poor quality of life, only a handful of cities have successfully adopted congestion pricing [^2] At the time of writing, those are Singapore, London, Stockholm, Milan, Gothenburg, Oslo, Manhattan, and several other smaller cities and old town centers.. A typical congestion pricing policy charges drivers entering the city center a time-varying toll aimed at mimicking the real-time optimal price that would keep the city free of congestion. The case of Manhattan, NYC is particularly interesting, as it resisted congestion pricing for almost 73 years despite being one of the most congested cities in the world. Namely, William Vickrey, widely regarded as the father of congestion pricing (Vickrey, 1969), originally conceptualized it for the NYC subway system in 1952, while the Manhattan policy finally came into effect in early 2025. The long debate featured several rejected proposals (e.g., in 2008 and 2015) as well as protests and lawsuits against the policy, due to concerns that it would hurt marginalized communities that rely on driving to Manhattan from the boroughs (Burkett, 2024), small business owners (Trapani, 2023), and others (John Dias, 2023). As a consequence of this controversy, the final policy was implemented in January 2025 at a peak toll price of \$9, instead of the originally proposed \$15, which was already a conservative approximation of the theoretically optimal price to eliminate congestion. Notice that, other than tolling entry into a whole city, there are other real-world implementations of congestion pricing on smaller scales, notably the highway express lanes in California and other US states that largely resemble our running example (cf. Figure 3). These so-called High Occupancy/Toll (HOT) lanes have faced less controversy than tolling a whole city, or even the whole highway, since users at least have access to the free general purpose lanes (Decorla-Souza, 2008). Nonetheless, practical limitations on how high the toll can be persist, and HOT lanes have been nicknamed “Lexus lanes” due to public perception that they are “built for the rich” (Patterson and Levinson, 2008).
@@ 2.3 Pillar 3: Embrace the dynamics
Example 3 demonstrates that repetition, coupled with dynamics that link the allocations across time, opens up new possibilities with respect to the single-shot and static setting. This is formalized by a recent but already seminal result by Jackson and Sonnenschein (2007). Stating this result requires extending the notation introduced in Section 2.1 to repeated and dynamic settings. A finite set $\mathcal{X}$$\subset{\mathbb{N}}$ is repeatedly allocated from over $T\in{\mathbb{N}}$ time-steps. In each time-step $t\in{\mathcal{T}}=\{1,\dots,T\}$, the private individual reward functions $r_{t,i}$ are drawn independently across time and users from a finite set $\tilde{\mathcal{R}}_{i}\subset{\mathcal{R}}_{i}$ according to publicly known probability distributions $\bm{\sigma_{i}}\in\Delta(\tilde{\mathcal{R}}_{i})$, with $\bm{\sigma}$ denoting the corresponding joint distribution. Let $f$ be a single-shot allocation rule as introduced in Section 2.1, and let $\bar{r}_{i}^{f}=\operatorname*{\mathbb{E}}_{\bm{r}\sim\bm{\sigma},x\sim f(\bm{r})}\left[r_{i}(x)\right]$ denote user $i$’s expected single-shot reward under $f$ (if all users report their rewards truthfully). Allocation rule $f$ is said to be ex-ante Pareto efficient if there is no dominating $f^{\prime}$ for which $\bar{r}_{i}^{f^{\prime}}\geq\bar{r}_{i}^{f}$ for all $i\in{\mathcal{N}}$ with strict inequality for at least one $i$. Consider the following linking mechanism aimed at implementing $f$ in a truth-revealing manner when repeated over $T$. Each user $i$ receives a budget to report each $r_{i}\in\tilde{\mathcal{R}}_{i}$ exactly $\sigma_{i}(r_{i})\,T$ times [^4] For the sake of simplicity we assume here that $\sigma_{i}(r_{i})\,T$ is integer for all $i\in{\mathcal{N}}$, $r_{i}\in\tilde{\mathcal{R}}_{i}$; c.f. Jackson and Sonnenschein (2007) for the general treatment., and the allocation in each time-step $t$ is chosen according to $f$ evaluated at the present reports $\bm{\tilde{r}_{t}}$. Let $\bm{\pi_{i}}=\left(\bm{\pi_{t,i}}\right)_{t\in{\mathcal{T}}}$ denote user $i$’s policy under this mechanism that specifies a probability distribution over its reports at each time-step (as a function of its true reward, the history of reports and allocations, etc.), and $\bm{\pi}=\left(\bm{\pi_{i}}\right)_{i\in{\mathcal{N}}}$ the policy profile of all users. When policy profile $\bm{\pi}$ is followed, each user experiences an expected average reward of $\bar{r}_{i}(\bm{\pi})=\cfrac{1}{T}\,\operatorname*{\mathbb{E}}\limits_{\begin{subarray}{c}\bm{\tilde{r}_{t}}\sim\bm{\pi_{t}}\\ x_{t}\sim f(\bm{\tilde{r}_{t}})\\ t\in{\mathcal{T}}\end{subarray}}\left[\sum\limits_{t\in{\mathcal{T}}}r_{t,i}(x_{t})\right]$. A policy profile $\bm{\pi^{\star}}$ is an $\epsilon$-approximate Bayesian Nash equilibrium if no user can unilaterally improve its expected average reward by more than $\epsilon$, i.e., if $\bar{r}_{i}\left(\bm{\pi_{i}^{\star}},\bm{\pi^{\star}_{-i}}\right)\geq\bar{r}_{i}\left(\bm{\pi_{i}},\bm{\pi^{\star}_{-i}}\right)-\epsilon$ for all users $i\in{\mathcal{N}}$ and policies $\bm{\pi_{i}}$. A policy profile in which all users report their rewards as truthfully as possible [^5] In the ex-post realization of rewards, it may not be possible to report them truthfully in every time-step under the linking mechanism, and the reports are as truthful as possible if they are truthful except for a minimal number of time-steps; cf. Jackson and Sonnenschein (2007) for the formal definition. is denoted by $\bm{\pi^{\textup{true}}}$.
@@
Theorem 2 essentially shows that dynamics enable escaping the static impossibility of jointly Pareto efficient, truthful, and non-dictatorial allocation rules, as established in Theorem 1. To make this claim rigorous, some subtleties should be pointed out between both theorems. The first subtlety is that Theorem 1 concerns ex-post Pareto efficiency while Theorem 2 concerns ex-ante Pareto efficiency; however, the ex-ante notion is stronger and leads to even less possibilities in the single-shot setting (Hylland, 1980, Theorem 2) [^6] In randomized allocations, “ex-ante” refers to expected quantities before randomization takes place, and “ex-post” to realized quantities after randomization. The only possibility to attain ex-ante Pareto efficiency and strategy-proofness in the single-shot setting is a (non-random) dictatorship.. The second subtlety is that Theorem 1 concerns strategy-proofness, i.e., truthful reporting is a dominant strategy; while Theorem 2 concerns so-called Bayesian Nash incentive compatibility, i.e., truthful reporting is an ($\epsilon$-approximate) Bayesian Nash equilibrium, which is weaker than strategy-proofness. However, it is generally believed, though difficult to formally prove, that relaxing strategy-proofness to Bayesian Nash incentive compatibility does not open significant possibilities in the single-shot setting. Some impossibility results for Bayesian Nash incentive compatibility have also been established recently (Ehlers et al., 2020; Kikuchi and Koriyama, 2024).
While Theorem 2 provides hope that trustworthy, fair and efficient dynamic resource allocation schemes are feasible, the linking mechanism used in the theorem is not well-suited for socio-technical control. To summarize the shortcomings of this mechanism in control-theoretic terms, it is ‘open-loop.’ Namely, it requires precise knowledge of the reward function distributions $\bm{\sigma_{i}}$ to issue the correct amount of tokens. Moreover, it is not suitable for allocations that are repeated indefinitely, as is the case in most socio-technical control contexts, since it would require issuing infinite tokens. As discussed in Section 4.5, these shortcomings are addressed by our proposed ‘closed-loop’ counter-part, that is, karma economies.